OpenStudy (anonymous):

Solve: 4(2x − 1) − 3 = 9(x − 4) Can you help to solve this by Pam

OpenStudy (anonymous):

Expand both sides, collect x-terms on the left, and numbers on the right and divide where you have to"\[4(2x-1)-3=9(x-4) \rightarrow 8x-7=9x-36\]\[8x-9x=-36-(-7)\rightarrow -x=-29\rightarrow x=29\]

OpenStudy (anonymous):

that would not be -29 just 29 right

OpenStudy (anonymous):

yeah 29

OpenStudy (anonymous):

you have to divide the second-last one by -1 on both sides to get positive x = positive 29

OpenStudy (anonymous):

Thank you